OpenStudy (anonymous):

When negative five is subtracted from a number the result is 10. Find the number.

OpenStudy (some.random.cool.kid):

welcome to openstudy !!!! :}

OpenStudy (anonymous):

hey

OpenStudy (some.random.cool.kid):

this would be 15 because x-5=10 x=10+5 x=15

OpenStudy (anonymous):

thnaks